i+j and j+k should be perpendicular vectors (by simple visualisation) then why is their dot product not 0 but 1?

Grade:12

5 Answers

Aryan
6 years ago
, while visiting flowers, they come in contact and adhere pollen grain and bring about pollination, which they visit another flower. ... Protein pellicle and stigmatic exudates ...
Aryan
6 years ago
, while visiting flowers, they come in contact and adhere pollen grain and bring about pollination, which they visit another flower. ... Protein
Aryan
6 years ago
Dkdmn, while visiting flowers, they come in contact and adhere pollen grain and bring about pollination, which they visit another flower. ... Protein pellicle and stigmatic exudates ...
Aryan
6 years ago
Krkr, while visiting flowers, they come in contact and adhere pollen grain and bring about pollination, which they visit another flower. ... Protein pellicle and stigmatic exudates ...
Kashinath Patekar
6 years ago
The dot product of the two unit vectors is not 0 but 1, since they are NOT perpendicular to each other. In fact they are inclined at an angle of 60° to each other.Explanation:-Let the two vectors be a=i+j and b=j+kLet the angle between two vectors be AThus, CosA = { |a.b| } / { |a|*|b| }|a.b|=1 and |a|=|b|=✓(1²+1²) = ✓2Therefore,CosA= 1/(✓2*✓2)CosA= 1/2Hence, A=60°